OpenStudy (ahammock):

I don't know how to solve for the number atoms for Carbon, Hydrogen, and Oxygen in this problem. Please help if you can. Calculate the number of molecules of the compound and the number of atoms of each of the elements in 2.50 moles of H2C2O4. The molecule total is 1.50e24 atoms C atoms H atoms O

OpenStudy (frostbite):

Hey @Ahammock and welcome to Openstudy... or what is left of it. First of: Remember the number of molecules in total must be rounded up as you get 1.505, which yields 1.51e24. Aside from the minor detail lets go to what matters: the amount of mole of each atom: Remember that mole is simply a multiplicative prefix like kilo, mega, nano and so on, meaning that if you know the amount of mole of a specie we can find the total amount of species (like you did for the whole molecule: For simplicity I will give you a way so you can solve similar reactions in the future: Lets look at the total chemical decomposition of \(\sf H_2C_2O_4\): \[\Large \sf H_2C_2O_4 \rightarrow 2~H + 2~C+4~O\] This reaction tells us: If you have 1 mole of \(\sf H_2C_2O_4\) then you have 2 mole H, 2 mole C and 4 mole O: The amount of moles for each element is therefore: Carbon: \(\Large n \left( \sf C \right)=2 \times n \left( \sf H_2 C_2 O_4 \right)\) Hydrogen: \(\Large n \left( \sf H \right)=2 \times n \left( \sf H_2 C_2 O_4 \right)\) Oxygen: \(\Large n \left( \sf O \right)=4 \times n \left( \sf H_2 C_2 O_4 \right)\) From here it is simply to multiply Avogadro's constant onto amount of substance in mole. Hope this can help you to finish your work.